gjoy:

Find the total area and volume for a cylinder with a radius of 1 unit and a height of 1 unit. The number of square units of the total area exceeds the number of cubic units in the volume by _____ units. π 2 π 3 π

AZ:

Total area or surface area of a cylinder can be found by: \(\large\text{SA}=2\pi rh+2\pi r^2\) The volume of a cylinder is \(\large \text{Volume} = \pi r^2h\)

AZ:

You're given radius = r = 1 height = h = 1 so put those numbers into your formula, what do you get? \(\large\text{SA}=2\pi rh+2\pi r^2\) \(\large\text{SA}=(2\times \pi \times 1 \times 1)+ (2 \times \pi \times 1^2)\) What is 2 * pi * 1 * 1 = ? just keep pi as pi, don't use 3.14 and similarly what is 2 * pi * 1^2 = and then add those two numbers you get, and that's your total area or surface area

AZ:

For volume, we use the same radius = 1 and height = 1 \(\large \text{Volume} = \pi r^2h\) \(\large \text{Volume} = \pi \times (1)^2 \times 1\) What is pi * 1^2 * 1 remember 1^2 is the same thing as 1 * 1

AZ:

Now to answer your question, all you have to do is Area - Volume = ?